### v2

# Parallelizing Linear Transformers with the Delta Rule over Sequence Length

Abstract
Transformers with linear attention (i.e., linear transformers) and state-space models have recently been suggested as a viable linear-time alternative to transformers with softmax attention. However, these models still underperform transformers especially on tasks that require in-context retrieval. While more expressive variants of linear transformers which replace the additive outer-product update in linear transformers with the delta rule [DeltaNet; ] have been found to be more effective at associative recall, existing algorithms for training such models do not parallelize over sequence length and are thus inefficient to train on modern hardware. This work describes a hardware-efficient algorithm for training linear transformers with the delta rule, which exploits a memory-efficient representation for computing products of Householder matrices []. This algorithm allows us to scale up DeltaNet to standard language modeling settings. We train a 1.3B model for 100B tokens and find that it outperforms recent linear-time baselines such as Mamba [] and GLA [] in terms of perplexity and zero-shot performance on downstream tasks (including on tasks that focus on recall). We also experiment with two hybrid models which combine DeltaNet layers with (1) sliding-window attention layers every other layer or (2) two global attention layers, and find that these hybrid models outperform strong transformer baselines. The parallel DeltaNet layer is made available as part of the FlashLinearAttention library [, ]: https://github.com/sustcsonglin/flash-linear-attention
1 Introduction
The attention mechanism [, ] has been shown to be an important primitive for accurate sequence modeling. Attention is moreover efficient during training as it is rich in matrix multiplications and can thus take advantage of highly parallel processing capabilities and specialized accelerators on modern GPUs. However, the complexity of attention is quadratic in sequence length, and hence it is a fundamentally expensive primitive. And while recent techniques have made it possible to scale attention to longer sequences through hardware-aware restructuring of the intermediate computations [, , , ], these methods still require storing the key/value vectors of previous elements, and this “KV cache” can be unwieldy to manage for long sequences.
Linear attention transformers [] replace the exponential kernel in softmax attention with a dot-product over (possibly transformed) key and query vectors. This makes it possible to formulate linear attention as a linear RNN with matrix-valued hidden states, thus obviating the need for a KV cache and enabling constant-memory inference. While initial variants of linear attention generally underperformed softmax attention on language modeling, gated variants of linear attention which incorporate a data-dependent gating factor have recently been shown to be competitive against strong transformer baselines [, , , ]. These gated linear transformers, along with time-varying state space models such as Mamba [] (which can be reparameterized as a gated linear transformer), have been suggested as a potential alternative to ordinary transformers. However, despite the competitive language modeling performance, these models have been shown to underperform transformers on recall-intensive tasks [], which is important for many practical downstream tasks of interest (e.g., in retrieval-augmented generation []).
To enhance associative recall over long contexts, [] propose DeltaNet, a variant of a linear transformer which uses a delta rule-like update [] to retrieve and update a value vector that is associated with the current key. DeltaNet was found to be effective on synthetic tasks and small scale language modeling/machine translation. However, the original work used a sequential algorithm that did not parallelize across sequence length, thus resulting in hardware-inefficient training, and it has not been clear how to scale DeltaNet to larger models and datasets.
This work describes a hardware-efficient training algorithm for DeltaNets which parallelizes the forward/backward passes across sequence length. We reparameterize the DeltaNet as a matrix-valued RNN whose recurrence is given by a generalized Householder transformation. This reparameterization enables the use of the memory-efficient WY representation [] for products of Householder matrices, eliminating the need to materialize the matrix-sized hidden states (which would incur high I/O cost). The memory-efficient representation makes it possible to straightforwardly extend the chunkwise parallel strategy for training linear attention models [, , ] to the DeltaNet case. We scale DeltaNets to moderate-scale language modeling benchmarks (1.3B models trained on 100B tokens), where DeltaNet is found to obtain better language modeling and zero-shot downstream task performance than strong linear recurrent models such as Mamba [] and GLA []. For in-context retrieval and learning evaluation, we evaluate DeltaNet on synthetic benchmarks such as MQAR [], RegBench [], MAD [], as well as real data benchmarks tested in [], where it is again found to perform well against linear recurrent baselines. Finally, we experiment with a hybrid approach where we combine DeltaNet layers with sliding attention layers or global attention layers, and find that these hybrid models can improve upon ordinary transformers, as well as the pure DeltaNet transformer.
2 Background
2.1 Linear Transformer: Transformers with Linear Attention
Given a sequence of $d$-dimensional input vectors ${\bm{x}}_{1},\dots,{\bm{x}}_{L}$, transformers use the softmax attention mechanism to attend over the entire past,
${\bm{q}}_{t},\ {\bm{k}}_{t},\ {\bm{v}}_{t}$ $={\bm{W}}_{Q}{\bm{x}}_{t},{\bm{W}}_{K}{\bm{x}}_{t},{\bm{W}}_{V}{\bm{x}}_{t},$ ${\bm{o}}_{t}=\sum_{i=1}^{t}\frac{\exp({\bm{k}}_{i}^{{\cramped{{}^{\mathsf{T}}}}}{\bm{q}}_{t})}{\sum_{j=1}^{t}\exp({\bm{k}}_{j}^{\cramped{{}^{\mathsf{T}}}}{\bm{q}}_{t})}{\bm{v}}_{i},$
where ${\mathbf{W}}_{Q},{\mathbf{W}}_{K},{\mathbf{W}}_{V}\in\mathbb{R}^{d\times d},{\bm{q}}_{t},{\bm{k}}_{t},{\bm{v}}_{t},{\bm{o}}_{t}\in\mathbb{R}^{d}$. (Here we assume a single attention head). Linear attention [] replaces the exponential kernel $\exp({\bm{k}}_{i}^{\cramped{{}^{\mathsf{T}}}}{\bm{q}}_{t})$ with the dot-product $\phi({\bm{k}}_{i})^{\cramped{{}^{\mathsf{T}}}}\phi({\bm{q}}_{t})$ where $\phi:\mathbb{R}^{d}\to\mathbb{R}^{n}$ is a feature map. This makes it possible to rearrange computations to represent linear attention as a linear RNN with matrix-valued hidden states,
${\bm{o}}_{t}$ $=\sum_{i=1}^{t}\frac{\phi({\bm{k}}_{i})^{{\cramped{{}^{\mathsf{T}}}}}\phi({\bm{q}}_{t})}{\sum_{j=1}^{t}\phi({\bm{k}}_{j})^{{\cramped{{}^{\mathsf{T}}}}}\phi({\bm{q}}_{t})}{\bm{v}}_{i}=\frac{\left(\sum_{i=1}^{t}{\bm{v}}_{i}\phi({\bm{k}}_{i})^{\cramped{{}^{\mathsf{T}}}}\right)\phi({\bm{q}}_{t})}{\left(\sum_{j=1}^{t}\phi({\bm{k}}_{j})^{{\cramped{{}^{\mathsf{T}}}}}\right)\phi({\bm{q}}_{t})}=\frac{{\mathbf{S}}_{t}\phi({\bm{q}}_{t})}{\mathbf{z}^{\cramped{{}^{\mathsf{T}}}}_{t}\phi({\bm{q}}_{t})},$
where ${\mathbf{S}}_{t}=\sum_{i=1}^{t}{\bm{v}}_{i}\phi({\bm{k}}_{i})^{\cramped{{}^{\mathsf{T}}}}\in\mathbb{R}^{d\times n}$ and $\mathbf{z}_{t}=\sum_{i=1}^{t}\phi({\bm{k}}_{i})\in\mathbb{R}^{n}$. If we allow $n$ to go to infinity, linear attention can use feature maps associated with polynomial kernels to compute a polynomial approximation to the exponential kernel as a dot product, and can thus approximate softmax attention arbitrarily well []. The denominator $\mathbf{z}_{t}^{\cramped{{}^{\mathsf{T}}}}\phi({\bm{q}}_{t})\in\mathbb{R}$ can result in numerical instabilities [] and is removed in many recent works [, , , ]. It is also common to use the identity mapping for $\phi$ [, ], which results in the following simplified linear transformer: ${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$, ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}$.
Efficient training.
Let ${\mathbf{Q}},{\mathbf{K}},{\mathbf{V}}\in\mathbb{R}^{L\times d}$ be the stacked query, key, value vectors, e.g., ${\mathbf{Q}}_{i}={\bm{q}}_{i}$. We can then compute the output ${\mathbf{O}}\in\mathbb{R}^{L\times d}$ in parallel via ${\mathbf{O}}=\left({\mathbf{Q}}{\mathbf{K}}^{\cramped{{}^{\mathsf{T}}}}\odot{\mathbf{M}}_{L}\right){\mathbf{V}}$, where ${\mathbf{M}}_{L}\in\mathbb{R}^{L\times L}$ is the causal mask. This fully “parallel form” and the above “recurrent form” have different FLOPs and memory cost tradeoffs. The parallel form takes $O(L^{2}d+Ld^{2})$ and thus requires more FLOPs than the recurrent form, which takes $O(Ld^{2})$. However, the parallel form can often be faster in practice for moderate-length sequences as it can be fully parallelized and done in $O(1)$ steps. This sequence-level parallellism also enables high GPU occupancy. While the recurrent form requires fewer FLOPs ($O(Ld^{2})$), the elementwise operations involved in recurrence have low arithmetic intensity, unlike the matmul operations in the parallel form which can make use of specialized compute units (e.g., tensor cores). Thus the recurrent form is often slower in practice. [^1] It is possible in theory to use parallel scan [] to parallelize the recurrent form, which would enable the computations to be performed in $O(\log L)$ steps and $O(Ld^{2})$ FLOPs. However, this approach requires materializing the 2D hidden state for each time step, which would incur significant memory I/O cost unless the hidden state size is small enough that the materialization can happen in faster memory (which is the strategy adopted by Mamba []).
Chunkwise parallel form.
The chunkwise parallel form [, ] strikes a balance between the parallel and recurrent forms, allowing for fewer FLOPs than the parallel form and more sequence-level parallelism than the recurrent form. Concretely, suppose the query/key/value vectors are split into $\frac{L}{C}$ chunks where each chunk is of length $C$. Let ${\mathbf{Q}}_{[t]}\in\mathbb{R}^{Cd}$ be all the query vectors for chunk $t$, and let ${\bm{q}}_{[t]}^{i}={\bm{q}}_{tC+i}$ be the $i$-th query vector within the $t$’th chunk; the key/value chunks are defined similarly. Note that $t\in[0,L/C)$, $i\in[1,C]$. The state matrices are also re-indexed such that ${\mathbf{S}}_{[t]}^{i}={\mathbf{S}}_{tC+i}$, and we additionally define ${\mathbf{S}}_{[t]}^{0}={\mathbf{S}}_{[t-1]}^{C}$, i.e., the initial state of a chunk is the last state of the previous chunk. We can then obtain the following identity for the hidden state and output vector for the $r$-th element within the $t$-th chunk,
${\mathbf{S}}_{[t]}^{r}$ $={\mathbf{S}}_{[t]}^{0}+\sum_{i=1}^{r}{\bm{v}}_{[t]}^{i}{\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}},\qquad{\bm{o}}_{[t]}^{r}={\mathbf{S}}_{[t]}^{0}{\bm{q}}_{[t]}^{r}+\sum_{i=1}^{r}{\bm{v}}_{[t]}^{i}\left({\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}}{\bm{q}}_{[t]}^{r}\right).$
By further rewriting the intra-chunk computation based on the parallel form, we obtain following “chunkwise parallel form”,
${\mathbf{S}}_{[t+1]}$ $={\mathbf{S}}_{[t]}+{\mathbf{V}}_{[t+1]}^{\cramped{{}^{\mathsf{T}}}}{\mathbf{K}}_{[t+1]}\qquad\qquad\qquad\qquad\qquad\in\mathbb{R}^{d\times d},$ (1)
${\mathbf{O}}_{[t+1]}$ $={\mathbf{Q}}_{[t+1]}{\mathbf{S}}_{[t]}^{{\cramped{{}^{\mathsf{T}}}}}+\left({\mathbf{Q}}_{[t+1]}{\mathbf{K}}_{[t+1]}^{\cramped{{}^{\mathsf{T}}}}\odot{\mathbf{M}}_{C}\right){\mathbf{V}}_{[t+1]}\quad\in\mathbb{R}^{C\times d},$ (2)
where we let ${\mathbf{S}}_{[t]}={\mathbf{S}}_{[t]}^{0}$ to reduce notational clutter. With this form, information is propagated chunk-to-chunk through ${\mathbf{S}}_{[t]}$, and the intra-chunk states ${\mathbf{S}}^{i}_{[t]}$ for $i\in[1,C]$ need not be materialized, thus saving memory.
The complexity of the chunkwise parallel form is $O(LCd+Ld^{2})$, and the number of steps (without parallel scan) is $O(\frac{L}{C})$. Hence, $C=L$ recovers the fully parallel form and $C=1$ recovers the recurrent form. The chunkwise parallel form allows us to interpolate between the two forms, in essence trading off the number of sequential computations against sequence-level parallelism. In practice $C$ is set to a small constant (usually 64 or 128), allowing for subquadratic training. This chunkwise form enables practical speed-ups against parallel-form-only softmax attention even on moderate-length sequences, as demonstrated by FlashLinearAttention [, ]
2.2 DeltaNet: Linear Transformers with the Delta Update Rule
The above linear transformer employs a simple linear recurrence: ${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$. This can be seen as additively updating the memory ${\mathbf{S}}_{t-1}$ with new key-value associations at each time step. However, a purely additive update rule makes it difficult to deallocate past key-value associations, eventually leading to key “collisions” when $L>d$, as pointed out by []. A model should ideally learn to remove less important key-value associations to make room for new ones, and this removal should depend on the interaction between the new key and the memory content.
DeltaNet uses the delta update rule [] to operationalize this mechanism. Specifically, it first retrieves the old value using the current key, ${\bm{v}}_{t}^{\text{old}}={\mathbf{S}}_{t-1}{\bm{k}}_{t}$. It then obtains a new value ${\bm{v}}_{t}^{\text{new}}$ by interpolating between the old value and the current value ${\bm{v}}_{t}$,
${\bm{v}}_{t}^{\text{new}}=\beta_{t}{\bm{v}}_{t}+\left(1-\beta_{t}\right){\bm{v}}_{t}^{\text{old}},$
where $\beta_{t}=\sigma({\mathbf{W}}_{\beta}{\bm{x}}_{t})\in(0,1)$. This new value is added to the memory, resulting in the following update:
${\mathbf{S}}_{t}$ $={\mathbf{S}}_{t-1}\underbrace{-{\bm{v}}_{t}^{\text{old}}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}}_{\text{remove}}\underbrace{+{\bm{v}}_{t}^{\text{new}}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}}_{\text{write}}.$
Here $\beta_{t}$ is a “writing strength”: when $\beta_{t}=1$, the old value is completely removed and ${\bm{v}}_{t}^{\text{new}}={\bm{v}}_{t}$; when $\beta_{t}=0$, the memory remains unmodified and we have ${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}$. The output computation is the same as vanilla linear attention, i.e., ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}$. The complexity of this recurrent form is the same as that of vanilla linear attention, i.e., $O(Ld^{2})$. This DeltaNet is a special case of fast weight programmers [], and [] and [] show that this type of linear transformer outperforms ordinary linear transformers on small-scale language modeling and synthetic retrieval tasks.
3 Parallelizing DeltaNet Across Sequence Length
In the same spirit as the chunkwise form of linear attention, we derive a chunkwise form for DeltaNet that enables hardware-efficient training through parallelizing across sequence length.
3.1 A Memory-efficient Reparameterization
We first observe that ${\mathbf{S}}_{t}$ admits a purely additive representation of the form ${\mathbf{S}}_{t}=\sum_{i=1}^{t}{\bm{u}}_{i}{\bm{k}}_{i}^{\cramped{{}^{\mathsf{T}}}}$ for ${\bm{u}}_{i},{\bm{k}}_{i}\in\mathbb{R}^{d}$, since we can simply set ${\bm{u}}_{i}={\bm{v}}_{i}^{\text{new}}-{\bm{v}}_{i}^{\text{old}}=\beta_{i}({\bm{v}}_{i}-{\bm{v}}_{i}^{\text{old}})$. Recall from §2.1 that simple linear attention has the form ${\mathbf{S}}_{t}=\sum_{i=1}^{t}{\bm{v}}_{i}{\bm{k}}_{i}^{\cramped{{}^{\mathsf{T}}}}$. Thus, DeltaNet simply replaces the value vector ${\bm{v}}_{i}$ in linear attention with the “pseudo” value vector ${\bm{u}}_{i}$. Once the ${\bm{u}}_{i}$’s have been constructed, the rest of computation can proceed as in ordinary linear attention, i.e., ${\mathbf{O}}=\left({\mathbf{Q}}{\mathbf{K}}^{\cramped{{}^{\mathsf{T}}}}\odot{\mathbf{M}}\ \right){\mathbf{U}}$ where ${\mathbf{U}}\in\mathbb{R}^{L\times d}$ is the row-wise concatenation of the ${\bm{u}}_{i}$ vectors.
However, computing ${\bm{u}}_{t}$ naïvely requires explicitly materializing ${\mathbf{S}}_{t-1}$ to compute ${\bm{v}}_{t}^{\text{old}}$, which would require $O(d^{2})$ memory. We now show that we can obtain the ${\bm{u}}_{t}$’s without explicitly materializing ${\mathbf{S}}_{t-1}$ in $O(d)$ memory. Our simple proof (by induction) relies on an application of the WY representation for products of Householder matrices []. The base case is clear since we have ${\mathbf{S}}_{1}=\beta_{1}{\bm{v}}_{1}{\bm{k}}_{1}^{\cramped{{}^{\mathsf{T}}}}$, so ${\bm{u}}_{1}=\beta_{1}{\bm{v}}_{1}$. For the inductive step, we first observe that the DeltaNet update is given by,
${\mathbf{S}}_{t}$ $={\mathbf{S}}_{t-1}-{\bm{v}}_{t}^{\text{old}}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}+{\bm{v}}_{t}^{\text{new}}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}={\mathbf{S}}_{t-1}-\beta_{t}\left({\mathbf{S}}_{t-1}{\bm{k}}_{t}\right){\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}+\beta_{t}{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}={\mathbf{S}}_{t-1}({\mathbf{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})+\beta_{t}{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}},$
which can be seen as applying a generalized Householder transformation (i.e., matmul with an identity plus rank-one matrix) to the previous state. The inductive step is then given by,
${\mathbf{S}}_{t}$ $={\mathbf{S}}_{t-1}({\mathbf{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})+\beta_{t}{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}=\sum_{i=1}^{t-1}{\bm{u}}_{i}{\bm{k}}_{i}^{\cramped{{}^{\mathsf{T}}}}+\underbrace{\left(\beta_{t}{\bm{v}}_{t}-\beta_{t}\sum_{i=1}^{t-1}{\bm{u}}_{i}\left({\bm{k}}_{i}^{{\cramped{{}^{\mathsf{T}}}}}{\bm{k}}_{t}\right)\right)}_{{\bm{u}}_{t}}{\bm{k}}_{t}^{{\cramped{{}^{\mathsf{T}}}}}=\sum_{i=1}^{t}{\bm{u}}_{i}{\bm{k}}_{i}^{\cramped{{}^{\mathsf{T}}}}.$
Note that ${\bm{u}}_{t}$ does not require materializing any of the hidden states and requires $O(d)$ memory to compute, thus completing the proof. While we have successfully avoided materializing the ${\mathbf{S}}_{t}$’s, computing the ${\bm{u}}_{t}$’s for all $L$ (i.e., ${\mathbf{U}}$) takes $O(L^{2}d)$ and moreover cannot be fully parallelized, unlike in linear attention where we can calculate all the value vectors ${\mathbf{V}}$ in parallel in $O(1)$ steps. We thus seek more a efficient chunkwise strategy for parallelizing DeltaNet, which similarly exploits memory-efficient representations of products of Householder matrices.
3.2 Chunkwise Parallel Form for DeltaNet
To derive the chunkwise parallel form, we first unroll the recurrence,
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}({\mathbf{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})+\beta_{t}{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}=\sum_{i=1}^{t}\beta_{i}({\bm{v}}_{i}{\bm{k}}_{i}^{\cramped{{}^{\mathsf{T}}}})\left(\prod_{j=i+1}^{t}({\mathbf{I}}-\beta_{j}{\bm{k}}_{j}{\bm{k}}_{j}^{\cramped{{}^{\mathsf{T}}}})\right).$
We then define the following variables: ${\mathbf{P}}_{i}^{j}=\prod_{t=i}^{j}({\mathbf{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})\in\mathbb{R}^{d\times d}$, ${\mathbf{H}}_{i}^{j}=\sum_{t=i}^{j}\beta_{t}({\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}){\mathbf{P}}_{t+1}^{j}\in\mathbb{R}^{d\times d}$, where we let ${\mathbf{P}}_{i}^{j}={\mathbf{I}}$ whenever $i>j$. Intuitively, ${\mathbf{P}}_{i}^{j}$ is the “decay factor” to be applied to ${\mathbf{S}}_{i}$ for obtaining ${\mathbf{S}}_{j}$, and ${\mathbf{H}}_{i}^{j}$ represents the contributions to ${\mathbf{S}}_{j}$ starting from token $i$. (Hence ${\mathbf{S}}_{t}={\mathbf{H}}_{1}^{t}$). The chunkwise recurrence can then be written as,
${\mathbf{S}}_{[t]}^{r}={\mathbf{S}}_{[t]}^{0}{\mathbf{P}}_{[t]}^{r}+{\mathbf{H}}_{[t]}^{r}$ (3)
where we define the chunkwise variables ${\mathbf{S}}_{[t]}^{i}={\mathbf{S}}_{tC+i}$, ${\mathbf{P}}_{[t]}^{r}={\mathbf{P}}_{tC+1}^{tC+r}$, ${\mathbf{H}}_{[t]}^{r}={\mathbf{H}}_{tC+1}^{tC+r}$. Here we have $\frac{L}{C}$ chunks of size $C$. The trick is to now efficiently represent the ${\mathbf{P}}_{[t]}^{r},{\mathbf{H}}_{[t]}^{r}\in\mathbb{R}^{d\times d}$ matrices using a similar approach described in §3.1, so that these matrices can be stored in $O(d)$ memory. This is given by,
${\mathbf{P}}_{[t]}^{r}$ $={\mathbf{I}}-\sum_{i=1}^{r}{\bm{w}}_{[t]}^{i}{\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}},\qquad{\mathbf{H}}_{[t]}^{r}=\sum_{t=1}^{r}{\bm{u}}_{[t]}^{i}{\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}}$ $\in\mathbb{R}^{d\times d}$ (4)
${\bm{w}}_{[t]}^{r}$ $=\beta_{[t]}^{r}\left({\bm{k}}_{[t]}^{r}-\sum_{i=1}^{r-1}\left({\bm{w}}_{[t]}^{i}({\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}}{\bm{k}}_{[t]}^{r})\right)\right)$ $\in\mathbb{R}^{d}$ (5)
${\bm{u}}_{[t]}^{r}$ $=\beta_{[t]}^{r}\left({\bm{v}}_{[t]}^{r}-\sum_{i=1}^{r-1}\left({\bm{u}}_{[t]}^{i}({\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}}{\bm{k}}_{[t]}^{r})\right)\right)$ $\in\mathbb{R}^{d}$ (6)
The derivations for the above can be found in the appendix. Subsequently, based on Eq. 3, we can obtain the chunk-level recurrence for hidden states and outputs as,
${\mathbf{S}}_{[t]}^{r}$ $={\mathbf{S}}_{[t]}^{0}-\left({\mathbf{S}}_{[t]}^{0}\sum_{i=1}^{r}{\bm{w}}_{[t]}^{i}{\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}}\right)+\sum_{i=1}^{r}{\bm{u}}_{[t]}^{i}{\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}}={\mathbf{S}}_{[t]}^{0}+\sum_{i=1}^{r}\left({\bm{u}}_{[t]}^{i}-{\mathbf{S}}_{[t]}^{0}{\bm{w}}_{[t]}^{i}\right){\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}},$
${\bm{o}}_{[t]}^{r}$ $={\mathbf{S}}_{[t]}^{r}{\bm{q}}_{[t]}^{r}={\mathbf{S}}_{[t]}^{0}{\bm{q}}_{[t]}^{r}+\sum_{i=1}^{r}\left({\bm{u}}_{[t]}^{i}-{\mathbf{S}}_{[t]}^{0}{\bm{w}}_{[t]}^{i}\right)\left({\bm{k}}_{[t]}^{i{\cramped{{}^{\mathsf{T}}}}}{\bm{q}}_{[t]}^{i}\right).$
Letting ${\mathbf{S}}_{[t]}={\mathbf{S}}_{[t]}^{0}$, the above can be simplified to matrix notations similarly to Eq.1-2,
${\mathbf{S}}_{[t+1]}$ $={\mathbf{S}}_{[t]}+\left({\mathbf{U}}_{[t+1]}-{\mathbf{W}}_{[t+1]}{\mathbf{S}}_{[t]}^{{\cramped{{}^{\mathsf{T}}}}}\right)^{\cramped{{}^{\mathsf{T}}}}{\mathbf{K}}_{[t+1]}$ (7)
${\mathbf{O}}_{[t+1]}$ $={\mathbf{Q}}_{[t+1]}{\mathbf{S}}_{[t]}^{\cramped{{}^{\mathsf{T}}}}+({\mathbf{Q}}_{[t+1]}{\mathbf{K}}_{[t+1]}^{{\cramped{{}^{\mathsf{T}}}}}\odot{\mathbf{M}})\left({\mathbf{U}}_{[t+1]}-{\mathbf{W}}_{[t+1]}{\mathbf{S}}_{[t]}^{\cramped{{}^{\mathsf{T}}}}\right)$ (8)
where $\square_{[t]}=\square_{[t]}^{1:C}\in\mathbb{R}^{C\times d}$ for $\square\in\{{\mathbf{Q}},{\mathbf{K}},{\mathbf{V}},{\mathbf{O}},{\mathbf{U}},{\mathbf{W}}\}$ defines the chunkwise matrices that are formed from stacking the ${\bm{q}}_{t},{\bm{k}}_{t},{\bm{v}}_{t},{\bm{o}}_{t},{\bm{u}}_{t},{\bm{w}}_{t}$ vectors.
Training speed.
The sequential computations needed for constructing the chunkwise matrices ${\mathbf{U}}_{[t]}$’s and ${\mathbf{W}}_{[t]}$’s takes $O(Ld)$ work but can be parallelized across chunks, reducing the number of sequential steps from $L$ to $C$. Once these are obtained, the chunkwise hidden state updates (Eq. 7) and the output computations (Eq. 8) are largely the same as in linear attention, and thus take $O(LCd+Ld^{2})$ work and $O(\frac{L}{C})$ sequential steps.
$L$ $d_{\text{head}}$ Speed-up
2048 64 5.5x
4096 64 7.6x
8192 64 11.5x
2048 128 8.9x
4096 128 13.2x
2048 256 13.7x
Table 1: Speed comparison between the recurrent and chunkwise algorithms for DeltaNet on a single H100 GPU.
In order to save even more memory, we discard the chunk-level hidden states ${\mathbf{S}}_{[t]}$’s (which would take $O(\frac{L}{C}d^{2})$ space) after the forward pass and recompute them during the backward pass; the ${\mathbf{U}}$ and ${\mathbf{W}}$ matrices, which take up $O(Ld)$ space, are kept in memory.
We implement both the pure recurrent form and the chunkwise parallel form in Triton [] and show the speed-ups for various sequence lengths (L) and head dimensions ($d_{\text{head}}$) in Table 1, where the model dimension $d$ is 2048.222So far we have been assuming a single head ($d_{\text{head}}=d$) for easier exposition. In practice we use multiple heads where the head dimension $d_{\text{head}}$ is smaller than the model dimension $d$. In this case we have ${\mathbf{S}}_{t}\in\mathbb{R}^{d\times d_{\text{head}}}$. Note that our recurrent form kernel [^3] Our recurrent form kernel as well as the kernel of the original DeltaNet paper does not use parallel scan, as this would be impractical. See footnote 1. is already 2x faster than the CUDA recurrent form kernel provided in the original DeltaNet paper. [^4] https://github.com/IDSIA/recurrent-fwp/blob/master/algorithmic/fast_weight/fast_weight_cuda.cu. Our chunkwise algorithm enjoys greater speed-ups as the length $L$ and head dimension $d_{\text{head}}$ increases.
3.3 DeltaNet Transformer
We describe how the DeltaNet layer primitive is used to build up a transformer-like model using standard modules. We largely follow the LLaMA-architecture transformer [Transformer++, ] and just replace the self-attention layer with the DeltaNet layer. We also apply normalization before output projection to stable training [, ]. As the additional parameters for computing scalar $\beta_{t}$ terms are negligible, parameter allocation is roughly the same as in Transformer++: $4d^{2}$ for the DeltaNet layer and $8d^{2}$ for the SwiGLU FFN layer [].
Feature map and normalization.
[] originally follow [] and apply a “$\text{ELU}+1$[] to nonlineary transform the key/query vectors. We instead use the SiLU activation [], which was found to perform better (as reported by []). [] also normalize the key/query vectors with the $L_{1}$ norm. We use the $L_{2}$ norm instead. Normalization has two effects: ignoring $\beta_{t}$, both $L_{1}$- and $L_{2}$-normalization would ensure that ${\mathbf{I}}-{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$ has eigenvalues $\leq 1$, making training stable; moreover, ${\mathbf{I}}-{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$ is a projection matrix with $L_{2}$-normalization, and only erases information in one subspace while keeping the other $d-1$ subspace intact, which is beneficial for retaining information while still enabling some forgetting. Thus our key/query vectors are given by ${\bm{k}}_{t}=\frac{\text{SiLU}({\mathbf{W}}_{K}{\bm{x}}_{t})}{\|\text{SiLU}({\mathbf{W}}_{K}{\bm{x}}_{t})\|_{2}}$, ${\bm{q}}_{t}=\frac{\text{SiLU}({\mathbf{W}}_{Q}{\bm{x}}_{t})}{\|\text{SiLU}({\mathbf{W}}_{Q}{\bm{x}}_{t})\|_{2}}$. We ablate on these choices in our empirical study.
3.4 Hybrid Models
Following recent work on combining subquadratic token-mixing layers with existing neural network primitives [, , ], we also experiment with hybridizing DeltaNet models.
Convolutional layers.
Recent linear recurrent models typically incorporate a lightweight depthwise-separable convolution layer after the query/key/value projections [, , ]. This “short convolution” layer [] generalizes the shift SSM [], and is efficient in both number of parameters and computational cost. We also add a short convolution layer after the query/key/value projections.
Local sliding window and global attention.
Linear attention largely uses a content-based addressing mechanism [] and lacks positional information []. [] also argue that linear attention lacks the ability to perform precise local token shifts and comparisons, thus facing difficulties on retrieval-intensive tasks. Motivated by this, we experiment with two different hybrid architectures that incorporate softmax attention. We first explore local sliding window attention which attends over a sliding window and morever has been shown to significantly improve linear attention [, , , ]; we follow Griffin [] and interleave DeltaNet layers and sliding multi-query local attention (SMQA) layers. We also experiment with global attention, which has been found to be helpful [, ] even if only few of the recurrent layers are replaced with full global attention []; we follow [] to replace only two layers with global attention: the second layer and the $(\frac{N}{2}+1)$-th layer, where $N$ is total number of layers.
4 Empirical Study
We compare the DeltaNet against strong baselines in both synthetic and real-world language modeling settings. Our main baselines include: LLaMA-architecture Transformer++ []; RetNet [], a linear attention Transformer with non-data-dependent exponential decay and large head dimension; GLA [], a linear attention Transformer with data-dependent decay; and Mamba [], a selective state-space model with data-dependent decay. The number of parameters is kept (roughly) the same between all models.
4.1 Synthetic Benchmarks
641282565120255075100Model dimensionAccuracy (%)Sequence Length: 512, Key-Value Pairs: 64DeltaNetMambaGLARetNetRWKV4Hyena
Figure 1: Accuracy (%) on MQAR.
We evaluate on three synthetic benchmarks: Multi-query associative recall [MQAR; ], in-context language learning [RegBench; ], and Mechanistic Architecture Design [MAD; ].
MQAR evaluates language models’ ability to (in-context) recall information within a context when faced with multiple recall queries. [] show that the accuracy of MQAR has a strong correlation with the language modeling performance. We use []’s training setting and for DeltaNet we use 2 heads. We do not use convolutions for these experiments. Table 1 shows that DeltaNet performs perfectly (even without convolution) in the hardest setting of MQAR [], outperforming Mamba (which uses convolution) in low dimension settings.
[…]
Figure 2: Accuracy (%) on RegBench.
We next consider RegBench [], which is a synthetic dataset designed to assess in-context learning capability of different model architectures. Each input sequence in this benchmark consists of 10 to 20 strings drawn from a probabilistic finite automaton (PFA), where each example is separated by a special separator token. Each sequence is designed to be sampled from a distinct language defined by the PFA so that a model needs to infer the underlying language from the context on the fly. During testing, a model is evaluated on predicting the next token of testing sequences generated from held-out PFAs. We follow [] and perform a hyperparameter sweep for each model architecture and use the best result. Here again we find that DeltaNet performs strongly compared to baselines.
Finally, we consider the MAD benchmark [], a suite of synthetic token manipulation tasks designed to probe capabilities of model architectures. The results are shown in Table 2. Compared with other architectures, including MHA, DeltaNet is better at recalling tasks, especially on Fuzzy Recall, although it struggles on the “Memorize” task.
Model Compress Fuzzy Recall In-Context Recall Memorize Noisy Recall Selective Copy Average
Transformer 51.6 29.8 94.1 85.2 86.8 99.6 74.5
Hyena [] 45.2 7.9 81.7 89.5 78.8 93.1 66.0
Multihead Hyena [] 44.8 14.4 99.0 89.4 98.6 93.0 73.2
Mamba [] 52.7 6.7 90.4 89.5 90.1 86.3 69.3
GLA [] 38.8 6.9 80.8 63.3 81.6 88.6 60.0
DeltaNet 42.2 35.7 100 52.8 100 100 71.8
Table 2: Results on MAD. Results other than DeltaNet are directly borrowed from []. (Multihead) Hyena, DeltaNet and Mamba make use of convolutions, whereas GLA does not.
4.2 Language Modeling
Experimental setup.
Following prior work [, ], we evaluate on Wikitext perplexity and zero-shot common sense reasoning tasks, including LAMBADA [LMB.; ], PiQA [], HellaSwag [Hella.; ], WinoGrande [Wino.; ], ARC-easy (ARC-e) and ARC-challenge (Arc-c) []. Following [], we also evaluate the models real-world recall-intensive tasks, including FDA [], SWDE [], and SQUAD []. Both SWDE and FDA focus on extracting structured information: SWDE from raw HTML to identify semi-structured relationships, and FDA from PDFs to retrieve key-value pairs. SQUAD evaluates language models on reading comprehension by providing a text passage and a related question.
Hyperparameters.
We train all models from scratch in two configurations: 340M and 1.3B parameters. Each model uses AdamW for optimization, with a peak learning rate of 3e-4. The 340M models are trained using 15 billion tokens and a batch size of 0.5M tokens, while the 1.3B models are trained with 100 billion tokens and a batch size of 2M tokens. We use a cosine learning rate schedule, starting with a warm-up phase of 0.5 billion tokens for the 340M models and 1 billion tokens for the 1.3B models. Both configurations have initial and final learning rates set at 3e-5. We apply a weight decay of 0.01 and use gradient clipping at a maximum of 1.0. The head dimension of DeltaNet is set to 128, and the kernel size for convolution layers is set at 4.
Model Wiki. LMB. LMB. PIQA Hella. Wino. ARC-e ARC-c Avg. SWDE SQUAD FDA
ppl $\downarrow$ ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc_n $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc_n $\uparrow$ cont. $\uparrow$ cont. $\uparrow$ cont. $\uparrow$
340M params / 15B tokens
Transformer++ 28.39 42.69 31.0 63.3 34.0 50.4 44.5 24.2 41.2 42.2 22.1 21.4
RetNet (w/o. conv) 32.33 49.19 28.6 63.5 33.5 52.5 44.5 23.4 41.0 13.3 27.6 2.9
Mamba (w. conv) 28.39 39.66 30.6 65.0 35.4 50.1 46.3 23.6 41.8 12.4 23.0 2.1
GLA (w/o. conv) 28.65 43.35 30.3 64.8 34.5 51.4 45.1 22.7 41.5 18.6 27.2 8.1
            (w. conv) 29.47 45.53 31.3 65.1 33.8 51.6 44.4 24.6 41.8 24.0 24.7 7.3
DeltaNet (w/o. conv) 29.08 50.87 30.0 63.6 33.6 51.7 46.0 23.0 41.3 24.6 26.9 4.5
DeltaNet (w. conv) 28.24 37.37 32.1 64.8 34.3 52.2 45.8 23.5 42.1 26.4 28.9 12.8
       + Sliding Attn 27.06 38.17 33.4 64.0 35.3 50.9 45.9 23.2 42.1 39.3 32.5 18.8
       + Global Attn (2 layers) 27.51 35.04 33.5 64.0 34.5 51.7 46.0 23.3 42.1 42.9 32.1 23.1
1.3B params / 100B tokens
Transformer++ 16.85 13.44 48.9 70.8 49.6 53.6 56.0 26.5 50.9 66.6 31.5 27.4
RetNet (w/o. conv) 18.64 17.27 43.3 70.0 47.3 52.5 54.8 25.6 48.9 42.8 34.7 14.3
Mamba (w. conv) 17.06 13.89 46.2 72.2 40.1 54.1 59.0 28.2 50.0 41.4 35.2 6.2
GLA (w/o. conv) 17.22 14.47 46.9 71.8 49.8 53.9 57.2 26.6 51.0 50.6 42.6 19.9
            (w. conv) 17.25 14.92 46.2 70.6 49.9 53.0 55.3 27.0 50.4 52.4 37.4 22.3
DeltaNet (w. conv) 16.87 12.21 48.9 71.2 50.2 53.6 57.2 28.3 51.6 49.5 37.4 17.2
       + Sliding Attn 16.56 11.74 49.2 71.8 51.1 52.8 58.9 28.8 52.1 53.3 43.3 22.3
       + Global Attn (2 layers) 16.55 12.40 48.8 70.8 50.7 54.2 58.4 28.1 51.8 71.0 43.0 29.8
DeltaNet Ablations (340M)
       w. $L_{1}$-norm & 1+ELU 31.12 55.96 26.3 63.9 33.0 50.9 44.3 21.8 40.1 14.5 23.9 6.2
       w. $L_{2}$-norm & 1+ELU 28.03 37.62 32.2 65.7 34.7 51.8 45.4 22.5 42.1 23.8 28.6 13.1
       w. $L_{2}$-norm & ReLU 28.75 43.53 30.2 64.0 33.9 48.9 45.6 22.8 40.9 27.2 26.7 9.0
Table 3: Main language modeling results against Transformer++, RetNet [], Mamba [], and GLA []. All models are trained on the same subset of the SlimPajama dataset with the Mistral tokenizer. The Transformer++, RetNet, Mamba, GLA (w/o. conv) results are taking from []. For hybrid models, “Sliding Attn” interleaves a sliding window attention every other layer, and “Global Attn” uses full global attention on two layers. The 340M/1.3B models are trained for 15B/100B tokens respectively. All results are obtained through lm-evaluation-harness [].
Results.
Our main language modeling results are shown in Table 3. In the original works, Mamba uses convolutions by default, while GLA does not. For fair comparison, we retrain the GLA with convolution, and also train DeltaNet without convolution. For the 1.3B setting we only train the DeltaNet with convolution due to limited compute resources.
In general we find that DeltaNet outperforms the strong Mamba/GLA baselines in terms of both perplexity and downstream task performance. Improvements upon Mamba are particularly pronounced for the recall intensive tasks (i.e., SWDE, SQUAD, FDA). We also confirm the benefits of hybrid architectures [, ]. Both the sliding window attention hybrid and the global attention hybrid work well, outperforming the Transformer++ in terms of perplexity. The hybrid DeltaNet that just replaces two DeltaNet layers with global attention outperforms a Transformer++ even on recall-intensive benchmarks.
Ablation study.
In Table 3 (bottom) we ablate the choice of feature map and normalization. The original DeltaNet uses $L_{1}$-norm and the ELU feature map. We find that simply replacing the $L_{1}$-norm with the $L_{2}$-norm greatly increases performance. As for feature map, we experiment with $\{\mathtt{ReLU},\mathtt{1+ELU},\mathtt{SiLU}\}$ and find that $\mathtt{SiLU}$ performs the best.
4.2
${\mathbf{S}}_{t}\in\mathbb{R}^{d\times n}$${\bm{x}}_{t}\in\mathbb{R}^{d}$
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}\bullet{\mathbf{M}}_{t}+{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}},$
${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t},$
[…]
, ${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}$
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+{\bm{v}}_{t}\phi({\bm{k}}_{t})^{\cramped{{}^{\mathsf{T}}}}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}\phi({\bm{q}}_{t})$
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}+{\bm{v}}_{t}\phi({\bm{k}}_{t})^{\cramped{{}^{\mathsf{T}}}},\,\,\,{\bm{z}}_{t}={\bm{z}}_{t-1}+\phi({\bm{k}}_{t})$ ${\bm{o}}_{t}={\mathbf{S}}_{t}\phi({\bm{q}}_{t})/({\bm{z}}_{t}^{\cramped{{}^{\mathsf{T}}}}\phi({\bm{q}}_{t}))$
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}(\mathbf{I}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})+\beta_{t}{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}$
${\mathbf{S}}_{t}=g_{t}{\mathbf{S}}_{t-1}+(1-g_{t}){\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}},\,\,\,{\bm{z}}_{t}=g_{t}{\bm{z}}_{t-1}+(1-g_{t}){\bm{k}}_{t}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}/({\bm{z}}_{t}^{\cramped{{}^{\mathsf{T}}}}{\bm{q}}_{t})$
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}\odot\exp(-({\bm{\alpha}}\mathbf{1}^{\cramped{{}^{\mathsf{T}}}})\odot\exp({\bm{A}}))+{\bm{B}}\odot({\bm{v}}_{t}\mathbf{1}^{\cramped{{}^{\mathsf{T}}}})$ ${\bm{o}}_{t}=({\mathbf{S}}_{t}\odot{\bm{C}})\mathbf{1}+{\bm{d}}\odot{\bm{v}}_{t}$
, ${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}\odot({\bm{\beta}}_{t}{\bm{\alpha}}_{t}^{\cramped{{}^{\mathsf{T}}}})+{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}$
, ${\mathbf{S}}_{t}=\gamma{\mathbf{S}}_{t-1}+{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}$
${\mathbf{S}}_{t}={\mathbf{S}}_{t-1}\odot\exp(-({\bm{\alpha}}_{t}\mathbf{1}^{\cramped{{}^{\mathsf{T}}}})\odot\exp({\bm{A}}))+({\bm{\alpha}}_{t}\odot{\bm{v}}_{t}){\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}+{\bm{d}}\odot{\bm{v}}_{t}$
[…]
${\mathbf{S}}_{t}=f_{t}{\mathbf{S}}_{t-1}+i_{t}{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}},\,\,\,{\bm{z}}_{t}=f_{t}{\bm{z}}_{t-1}+i_{t}{\bm{k}}_{t}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}/\max\{1,|{\bm{z}}_{t}^{\cramped{{}^{\mathsf{T}}}}{\bm{q}}_{t}|\}$
${\mathbf{S}}_{t}=\gamma_{t}{\mathbf{S}}_{t-1}+{\bm{v}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}$ ${\bm{o}}_{t}={\mathbf{S}}_{t}{\bm{q}}_{t}$
[…]
$\{{\bm{x}}_{t}\}_{t=1}^{L}\mapsto\{{\bm{o}}_{t}\}_{t=1}^{L}$${\bm{o}}_{t}$${\bm{x}}_{j}$$j>t$, , , , , , , , , ,
4.24.2, , , , , , 23
${\mathbf{M}}_{t}=\gamma_{t}({\mathbf{I}}-{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})$$\gamma_{t}\in(0,1)$55However we found the DeltaNet + local sliding-window attention hybrid to generalize well, which could provide an appealing middle ground.
${\mathbf{P}}_{n}={\mathbf{I}}-\sum_{t=1}^{n}{\bm{w}}_{t}{\bm{k}}_{t}^{{\cramped{{}^{\mathsf{T}}}}}$
${\mathbf{P}}_{n}$ $=\prod_{t=1}^{n}({\mathbf{I}}-\beta_{t}{\bm{k}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})$
$={\mathbf{P}}_{n-1}({\mathbf{I}}-\beta_{n}{\bm{k}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}})$
$=({\mathbf{I}}-\sum_{t=1}^{n-1}{\bm{w}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})({\mathbf{I}}-\beta_{n}{\bm{k}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}})$
$={\mathbf{I}}-\sum_{t=1}^{n-1}{\bm{w}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}-\beta_{n}{\bm{k}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}+(\sum_{t=1}^{n-1}{\bm{w}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}})\beta_{n}{\bm{k}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}$
$={\mathbf{I}}-\sum_{t=1}^{n-1}{\bm{w}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}-\underbrace{\left(\beta_{n}{\bm{k}}_{n}-\beta_{n}\sum_{t=1}^{n-1}\left({\bm{w}}_{t}({\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}{\bm{k}}_{n})\right)\right)}_{{\bm{w}}_{n}}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}$
$={\mathbf{I}}-\sum_{t=1}^{n}{\bm{w}}_{t}{\bm{k}}_{t}^{{\cramped{{}^{\mathsf{T}}}}}$
${\mathbf{S}}_{n}=\sum_{t=1}^{n}{\bm{u}}_{t}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}$
${\mathbf{S}}_{n}$ $={\mathbf{S}}_{n-1}({\mathbf{I}}-\beta_{n}{\bm{k}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}})+\beta_{n}{\bm{v}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}$
$=\left(\sum_{t=1}^{n-1}{\bm{u}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}\right)({\mathbf{I}}-\beta_{n}{\bm{k}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}})+\beta_{n}{\bm{v}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}$
$=\sum_{t=1}^{n-1}{\bm{u}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}-\left(\sum_{t=1}^{n-1}{\bm{u}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}\right)\beta_{n}{\bm{k}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}+\beta_{n}{\bm{v}}_{n}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}$
$=\sum_{t=1}^{n-1}{\bm{u}}_{t}{\bm{k}}_{t}^{\cramped{{}^{\mathsf{T}}}}+\underbrace{\left(\beta_{n}{\bm{v}}_{n}-\beta_{n}\sum_{t=1}^{n-1}{\bm{u}}_{t}\left({\bm{k}}_{t}^{{\cramped{{}^{\mathsf{T}}}}}{\bm{k}}_{n}\right)\right)}_{{\bm{u}}_{n}}{\bm{k}}_{n}^{{\cramped{{}^{\mathsf{T}}}}}$
$=\sum_{t=1}^{n}{\bm{u}}_{t}{\bm{k}}_{n}^{\cramped{{}^{\mathsf{T}}}}$